OpenStudy (anonymous):

Solve the following 3 equations please

OpenStudy (anonymous):

1. \[\frac{ k }{ 11 }=8\] 2. \[-\frac{ 13 }{ 3 } = \frac{ n }{ 3 }\] 3. \[\frac{ 8 }{ 9 } = \frac{ a }{ 18 }\] Thanks :)

OpenStudy (anonymous):